OpenStudy (anonymous):

Okay, so I'm looking for the cosine of a triangle with the following values. I have the answer of 0.4815, but the angle translates to 61.2. Can someone explain why this is so? Values: a = 8 , b = 3 , c = 9. The "proper" formula used to solve this is cosA =b^2 + c^2 - a^2 / 2b(c) .

OpenStudy (anonymous):

I understand how to solve for the answer, my issue is when to get from the decimal response to a legitimate angle.

OpenStudy (anonymous):

That is, from 0.4815 to 61.2

OpenStudy (perl):

I dont see how you got 61.2

OpenStudy (anonymous):

This is a lesson that they've provided for me, and they're somehow telling me the decimal can change to that.

OpenStudy (perl):

oh wait :) one moment

OpenStudy (perl):



OpenStudy (anonymous):

What about it?

OpenStudy (perl):

the answer is 61.2 degrees, i get the same

OpenStudy (perl):

1.068 radians

OpenStudy (perl):

here is a law of cosines calculator https://www.google.com/search?q=law+of+cosines+calculator&ie=utf-8&oe=utf-8

OpenStudy (anonymous):

But I'd like to know how they were able to convert it, rather than just using a calculator. Unless, that is what they did..

OpenStudy (perl):



OpenStudy (amistre64):

cos(A) = .4815 you have to take the inverse cosine to determine A

OpenStudy (amistre64):

if your calculator is set to degrees, itll give you: cos^(-1) (.4815) = 61.23

OpenStudy (amistre64):

if its set to radians, then youll either have to set it to degrees, or adjust the results by 180/pi

OpenStudy (anonymous):

It didn't give me the desired value.

OpenStudy (amistre64):

what did it give you?

OpenStudy (perl):

$$ \Large{ \cos A^o = 0.4185 \iff A^o = \cos^{-1}( 0.4185) } $$

OpenStudy (anonymous):

Ah, never mind. I was in radian mode. Whoops. You're right, thank you.

OpenStudy (amistre64):

yeah radians is 1.0684...

OpenStudy (anonymous):

¯\_(ツ)_/¯

OpenStudy (anonymous):

Alright, thanks for all the help.

OpenStudy (amistre64):

good luck